### v3

# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

Abstract
Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and text-to-SQL generation. While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist. Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and mathematics challenges. In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence). To assist LLMs in generating high quality responses, we present two prompting techniques: Miniature & Mull and Explain & Compare. The former technique is used to evaluate the semantic equivalence in which it asks LLMs to execute a query on a simple database instance and then explore if a counterexample exists by modifying the database. The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences. Our experiments demonstrate that using our techniques, LLMs are promising tools to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and have better alignment with human preference on text-to-SQL generations than execution accuracy.
1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale . Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024a). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics. However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance. We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al. (2020a). These methods assess query semantic equivalence by generating numerous fuzzy database instances, leading to significant computational and storage demands. Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al. (2020a).
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detecting implicit contents Huang et al. (2023); Zhang et al. (2023c), and solving complex logical problems Feng et al. (2023); Yao et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
  • Can LLMs help data engineers in identifying SQL semantic equivalence?
  • Should LLMs be used to evaluate text-to-SQL generation?
2 Background
2.1 Motivations
We first introduce a query rewriting error that happened in the real-world system, known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
1Q1:
2SELECT Parts.pnum FROM Parts
3WHERE Parts.qoh = (SELECT COUNT(Supply.shipdate)
4 FROM Supply
5 WHERE Supply.pnum = Parts.pnum
6 AND Supply.shipdate < 10);
7
8Q2:
9WITH Temp AS
10 SELECT Supply.pnum, COUNT(Supply.shipdate) AS ct
11 FROM Supply
12 WHERE Supply.shipdate < 10
13 GROUP BY Supply.pnum
14SELECT Parts.pnum FROM Parts,Temp
15WHERE Parts.qoh = Temp.ct
16AND Parts.pnum = Temp.pnum;
One may think these two queries are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate is less than 10; however they are different in how they handle the subquery. Suppose that there is a record $R$ in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table. Then, Q1 will include record $R$ since R.qoh is zero and selecting the count for R.pnum from the Supply table will also return zero. On the other hand, Q2 will not include record $R$ because the TEMP result will not contain R.pnum.
Hence, real-world data systems need tools to avoid such pitfalls. While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly. For instance, when we present the COUNT BUG to ChatGPT, ChatGPT is able to correctly identify that these two SQL queries are not equivalent.
Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al. (2024b), and as a result, they can be interpreted and answered in multiple ways Min et al. (2020). From the user’s perspective, different SQL queries that convey the same logical answer to a question may all be considered equivalent. However, the widely used execution match metric, which expects a single correct SQL query, does not capture these ambiguities. Consequently, we are interested in exploring whether LLMs can be utilized to assess the quality of text-to-SQL generation.
2.1.1 Two Notions of SQL Equivalence
Motivated from the query optimizers in data systems and evaluation metrics for text-to-SQL, we define two notions of SQL equivalence.
Definition 2.1 (Semantic Equivalence).
Two SQL queries Q1 and Q2 are said to be semantically equivalent if and only if there does not exist a database instance such that Q1 and Q2 return different outputs.
Definition 2.2 (Relaxed Equivalence).
Two SQL queries Q1 and Q2 are said to be relaxed equivalent if one may edit a trivial portion of the SQL queries while preserving the same logic to make Q1 and Q2 semantically equivalent.
Figure 1: Demonstration of Counter Example, Miniature & MULL, and Explain % Compare prompts. The used SQL queries are not semantically equivalent but are relaxed equivalent.
2.2 Why LLM-SQL-Solver
First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases. Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub optimal performance. We believe that, with LLM-SQL-Solver, query optimizers can explore the vast universe of equivalent rewritten queries and then find the query plan with near optimal performance.
For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from a user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents). Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al. (2018). Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database. This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information. As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases and without accessing any stored data. Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether LLM-SQL-Solver (relaxed notion) can be used as a better metric for text-to-SQL benchmarks.
3 Proposed Method
Motivated by recent methods Deng et al. (2023); Yao et al. (2023); Wei et al. (2022), we present our prompting techniques to obtain high quality judgments from Open-AI GPT models. Our prompting technique adheres to key aspects widely adopted in NL2SQL works Li et al. (2023); Khatry et al. (2023); Liu et al. (2023); Taori et al. (2023b): 1) Clear Layout: We provide the instruction at the beginning, followed by the two SQL queries that need to be compared, and end by providing the database schema. 2) Clear Context: We provide the database schema and explicitly included the primary and foreign key constraints, since they are crucial context information for deciding SQL equivalence.
For semantic equivalence, we also added a short definition of SQL semantic equivalence (i.e., Note: Two SQL queries are semantically equivalent if and only if they yield identical results for all possible databases.). In addition, we first ask GPT to generate a counter example if there exists one and then decide if two SQL queries are equivalent. We call this prompt the Counter Example prompt. However, as shown in Figure 1, the counter example prompt leads to sub-optimal results, especially when GPT models only scratched the surface and observed the obvious (similarity between two SQL queries in structure and logic). This finding aligns with the prior works Wang et al. (2023); Zheng et al. (2023) which demonstrated that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limited reasoning capability.
To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask GPT to execute both queries on a simple database and then see if it can modify the database such that the two queries output different results (i.e., a counter example). As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent. This prompt follows the main ideas from the chain-of-thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1. We observe that LLMs are inclined to label two SQL queries as equivalent when they appear similar in structure. As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent.
4 Evaluations
We evaluate the capabilities of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries. In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models. The models are accessed via OpenAI API.
Hyper-parameter: We use zero demonstrations in our prompts. In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
4.1 Data Source
First, we compare our proposed Miniature&Mull prompt with the CounterExample prompt on i) Non-equivalent Queries Chu et al. (2017) and ii) Equivalent Queries Zhong . These two data sources include pairs of SQL queries that have been expertly labeled as semantically non-equivalent or equivalent.
  • Nonequivalent Queries Chu et al. (2017): Four challenging and one easy questions, including the COUNT BUG, one exam question from CSE344 at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
  • Equivalent Queries Zhong et al. (2020a): A hundred pairs of expert labeled equivalent queries.
Next, we showcase results on the alignment of our proposed LLM-SQL-Solver judgements with human preferences on semantic and relaxed notions using the Spider benchmark Yu et al. (2018). The Spider benchmark contains 1034 development questions covering 20 databases in different domains. The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extrahard (extra). We evaluate the judgments of LLM-SQL-Solver against execution accuracy (EX) based on human preferences. In all experiments, we use the open sourced EX evaluation code from GitHub  [^1] https://github.com/THU-BPM/ChatGPT-sql.
  • Chat Liu et al. (2023): Generate queries using ChatGPT with the zero-shot basic prompt.
  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting.
  • DAIL Gao et al. (2023): The current SOTA open sourced technique on the Spider benchmark. It uses the GPT-4 model with few-shot prompting and incorporates self-consistency Wang et al. (2022).
4.2 Judging Semantic Equivalence
Judge Prompt Noneq Eq
GPT 35 turbo CounterExample .8 .45
Miniature&Mull 1.0 $\uparrow$ .57 $\uparrow$
GPT 4 turbo CounterExample .6 .7
Miniature&Mull 1.0 $\uparrow$ .83 $\uparrow$
Table 1: Results in determining semantic non-equivalent and equivalent queries.
Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs. We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
Judge Data Equivalence/Prompt Easy Medium Hard Extra Acc
- Spider - 248 446 174 166 -
EX Chat - 221 331 101 72 70.1%
DIN - 228 382 132 101 81.5%
DAIL - 226 392 132 104 82.6%
GPT 3.5 turbo Chat Semantic/Miniature&Mull 210 293 84 48 61.4%
DIN Semantic/Miniature&Mull 207 307 96 76 66.3%
DAIL Semantic/Miniature&Mull 224 317 85 95 70.4%
[…]
DIN Relaxed/Explain&Compare 223 360 133 114 80.3%
DAIL Relaxed/Explain&Compare 218 371 140 119 82.0%
GPT 4 turbo Chat Semantic/ Miniature&Mull 199 244 78 55 55.7%
DIN Semantic/Miniature&Mull 209 308 99 73 66.6%
DAIL Semantic/Miniature&Mull 216 343 114 75 73.0%
[…]
DIN Relaxed/Explain&Compare 209 377 124 106 79.0%
DAIL Relaxed/Explain&Compare 220 382 141 110 82.5%
Table 2: A breakdown of the accuracy scores when using different metrics, i.e., EX and LLMs’ judgements with varying prompt techniques. The compared SQL queries are the gold SQL queries from Spider and the generated SQL queries from Chat, DIN, and DAIL.
As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the Counter Example prompt. When using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if the SQL queries are not equivalent), GPT 4 turbo failed on a user found bug originally posted on Hackers News lukaseder . This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries. Hence, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will). GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries. When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to generate a counter example), GPT 3.5 turbo and GPT 4 turbo both successfully identified all the mistakes. Our proposed Miniature & Mull prompt not only increased the accuracy of identifying non-equivalent SQL queries to 100% for both models but also increased the accuracy for identifying equivalent SQL queries by at least 12%. As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.
4.3 Judging NL2SQL Generation
There are two notions that one can use in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions. We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL). We observe that the relative ranking among Chat, DIN, and DAIL are preserved for both EX and LLM judgements. Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy. Since DAIL applied the self-consistency techniques (resource-intensive Gao et al. (2023)), we expect DAIL to be more likely to produce the best answers. As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
4.3.1 Agreement between LLMs’ Judgements and Execution Accuracy
For a majority of the questions, LLMs’ judgements and EX agree with each other. In particular, as shown in, Table 3, agreements between LLMs and EX for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and EX for the relaxed equivalence notion have at least 82.4%. We believe this is an encouraging signal that implies the good quality of the Spider benchmark and LLMs’ judgements. The Spider benchmark contains high quality data points (1000 human hours spent Yu et al. (2018)). Creating these high quality databases, which is necessary for good EX, are expensive and does not scale. Assuming the underlying database is well designed and the EX is highly accurate, then a high agreement between EX and LLMs are obviously desirable.
Judge Equiv Data Agreement
GPT 3.5 turbo Semantic Chat 69.6%
DIN 72.6%
DAIL 74.6%
[…]
DIN 78.9%
DAIL 84.6%
GPT 3.5 turbo Relaxed Chat 83.8%
DIN 88.4%
DAIL 90.0%
[…]
DIN 82.4%
DAIL 86.1%
Table 3: We compare the agreement between LLMs’ judgements and execution matches.
4.3.2 Human Evaluation on Samples
Although, the agreements between execution accuracy and LLM’s judgments are high, there are still lots of questions with disagreements. In this section, we provide a detailed analysis on the sampled questions from the set of questions where EX and LLM’s judgement differs. Since there are hundreds of such questions (different results between EX and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and EX disagree.
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. Three of the authors manually checked the equivalence of the sampled questions where LLMs and EX disagree. In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective). To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries.
Judge Equiv Data Prefer
GPT 3.5 turbo Semantic Chat 2/10 $\downarrow$
DIN 4/10 $\downarrow$
DAIL 5/10 -
[…]
DIN 6/10 $\uparrow$
DAIL 5/10 -
GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
DIN 5/10 $\uparrow$
DAIL 6/10 $\uparrow$
[…]
DIN 6/10 $\uparrow$
DAIL 7/10 $\uparrow$
Table 4: We compare the human preference between LLM judge and EX on 10 randomly sampled questions from different data sources.
Semantic Equivalence
When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than EX, but GPT 4 turbo results in better alignments. When EX detects discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent. There are a few rare cases that we observed in which the SQL queries are semantically equivalent but EX determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by EX.
Figure 2: The rates of LLM correctly executed the SQL queries on its own provided simple database.
In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten. Diving deeper into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs contain only subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice these differences; Hallucination. GPT 3.5 turbo may return wrong results when executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo performs worse compared to GPT 4 turbo.
When GPT 4 turbo is used as the judge, it constantly produced more accurate or comparable results to EX, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. EX often failed to capture this difference. Although GPT 4 turbo’s judgment is much better, GPT 4 turbo may still miss some minor differences such as similar but different column names (such as song_name v.s. name), overlook the fact that extra columns are selected, and also sometimes hallucinates. As shown in Figure 2, we explored how well do LLMs execute and reason about the SQL queries on their own generated simple database. GPT 4 turbo has much higher correct operation rates which leads to much better responses compared with the weaker model. One common phenomena that lead to false answers in EX is due to the return of the empty tuple. When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are. Another interesting observation raised from a few comparisons is about the case sensitivity. OpenAI GPT models consider databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default. However, many popular database are by default case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013). Therefore, we operate under the assumption that the underlying database instance is case-insensitive when evaluating semantic equivalence.
Relaxed Equivalence. When using the Explain & Compare prompt for determining the relaxed equivalence, LLMs are consistently better aligned with human judgments than EX, as shown in Table 4. The relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinions of users. The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
The main reasons for why LLMs’ judgements are better aligned with human preference than EX is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value. Since EX strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. One interesting observation we found is about null values. GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that GPT models consider that null values can be inserted into a column with primary key constraint. In fact, the SQLite databases, by default, allow null values to be inserted to primary key column. However, in many database primary key constraints explicitly states that null values are not allowed Oracle . Based on the relaxed equivalence definition, we consider records in the primary columns to be not null. Although LLMs correctly consider that having null values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
5 Discussion and Limitation
In this discussion section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver. During the SQL rewrite phase of the query optimization engine, the input SQL query is transformed into a new SQL query that can be optimized more easily. The key challenge in query rewriting is to determine if the input query and the rewritten query are indeed semantic equivalent. The LLM-SQL-Solver can be directly used as a tool, during the rewriting phase, to check semantic equivalence and ensure bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur in the data system. Furthermore, the integration of the LLM-SQL-Solver as a text-to-SQL evaluation metrics using the relaxed equivalence notion could improve the overall quality of user experience and enable more human-like capabilities.
Concurrent with our research, the chain-of-table approach Wang et al. (2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt. However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
There are some limitations or areas of improvement to this work. First, we only evaluated the capabilities of the Large Language Models of OpenAI. There are other alternative models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of the GPT model family’s capabilities. In addition, our evaluations on Spider Benchmark were conducted over randomly sampled questions, in which we have to manually label the equivalence of the two different notions (semantic and relaxed). Since there are nearly ten thousand pairs of SQL queries, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL. Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
6 Conclusion
In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with our relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, our proposed LLM-SQL-Solver avoids accessing stored user data, and hence no sophisticated databases need to be constructed.