### v4

# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

Abstract
Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and text-to-SQL generation. While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist. Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and mathematics challenges. In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence). To assist LLMs in generating high quality responses, we present two prompting techniques: Miniature & Mull and Explain & Compare. The former technique is used to evaluate the semantic equivalence in which it asks LLMs to execute a query on a simple database instance and then explore if a counterexample exists by modifying the database. The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences. Our experiments demonstrate that using our techniques, LLMs are promising tools to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and have better alignment with human preference on text-to-SQL generations than execution accuracy.
1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale . Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024a). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018); Wang et al. (2022b) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics. However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance. We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al. (2020a). SQL queries in text-to-SQL benchmarks often include constructs such as EXISTS, UNION, UNION ALL, and database schema doesn’t no exclude NULL values, making off-the-shelf SQL provers (e.g., Cosette Chu et al. (2017) and WeTune Wang et al. (2022b)) unsuitable for direct use. Consequently, researchers have turned to fuzzy testing, which assesses query semantic equivalence by generating numerous fuzzy database instances. However, this approach imposes substantial computational and storage demands. Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al. (2020a). This substantial data requirement arises from a common issue in execution-based approaches: incorrectly classifying semantically inequivalent queries as equivalent due to the not having a counter-example included. For example, when the underlying database lacks tuples matching the predicate and join conditions, two semantically inequivalent queries can both produce an empty tuple as output. As a result, these queries will appear to match, regardless of their actual differences.
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgment Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning Saparov and He (2022), detecting implicit contents Huang et al. (2023); Zhang et al. (2023c), and solving complex logical problems Feng et al. (2023); Yao et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
Can LLMs help data engineers in identifying SQL semantic equivalence?
Should LLMs be used to evaluate text-to-SQL generation?
The paper is organized as follows: Section 2 motivates the need of LLMs in determining SQL equivalence and discusses the background information of two notions of SQL equivalence and LLMs basics. Section 3 introduces two prompting methods, Miniature and Mull and Explain and Compare, to improve LLMs’ ability to assess SQL queries semantic and pragmatic equivalence. Section 4 presents experimental results on assessing SQL equivalence using GPT-3.5-Turbo and GPT-4o respectively, applying our two proposed prompting techniques on three datasets from the Spider benchmark. Section 5 discusses the implications of the proposed LLM-SQL-Solver for SQL query validation and text-to-SQL evaluation, while also highlighting potential areas of improvement for our approach. Finally, Section 6 summarizes our contributions and concludes this work.
2 Background
2.1 Motivations
We first introduce a query rewriting error that happened in the real-world system, known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
One may think these two queries are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate is less than 10; however they are different in how they handle the subquery. Suppose that there is a record in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table. Then, Q1 will include record since R.qoh is zero and selecting the count for R.pnum from the Supply table will also return zero. On the other hand, Q2 will not include record because the TEMP result will not contain R.pnum.
Hence, real-world data systems need tools to avoid such pitfalls. While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly. For example, when we presented the COUNT BUG to ChatGPT, it was able to correctly identify that the two given SQL queries were not equivalent, effectively pointing out the underlying issue. Highlighting the potential of LLMs to complement traditional approaches by offering new insights.
Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al. (2024b), and as a result, they can be interpreted and answered in multiple ways Min et al. (2020). From the user’s perspective, different SQL queries that convey the same logical answer to a question may all be considered equivalent. However, the widely used execution match metric, which expects a single correct SQL query, does not capture these ambiguities. Consequently, we are interested in exploring whether LLMs can be utilized to assess the quality of text-to-SQL generation.
Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al. (2024b), allowing for multiple valid interpretations and corresponding answers Min et al. (2020). From a user’s perspective, different SQL queries that yield the same logical answer to a question are often considered equivalent. The primary objective of text-to-SQL systems is to align with human preferences and deliver outputs that meet user expectations. However, the commonly used execution match metric, which assumes a single correct SQL query, fails to account for these ambiguities. For example, consider a ground truth SQL query like "SELECT shipdate FROM Supply" and a predicted SQL query such as "SELECT shipdate, COUNT(shipdate) FROM Supply". Although these queries produce different output tables (making them execution unmatched), they convey the same underlying information. In such cases, a user might find the predicted query acceptable, even though it fails the strict execution match criterion. The limitations of execution match highlights the need for alternative approaches to evaluate the quality of text-to-SQL generation. On the other hand, large language models (LLMs) are often trained using Reinforcement Learning with Human Feedback (RLHF), which aligns their outputs more closely with human preferences and judgments Bai et al. (2022). We are particularly interested in exploring whether LLMs can provide a more nuanced assessment by identifying and validating diverse SQL queries that share the same intent, thereby achieving better alignment with human preferences.
2.1.1 Two Notions of SQL Equivalence
Motivated from the query optimizers in data systems and evaluation metrics for text-to-SQL, we define two notions of SQL equivalence.
Definition 2.1 (Semantic Equivalence Codd (1970)).
Two SQL queries Q1 and Q2 are said to be semantically equivalent if and only if there does not exist a database instance such that Q1 and Q2 return different outputs.
Definition 2.2 (Pragmatic Equivalence Baker (2018)).
Two SQL queries Q1 and Q2 are said to be pragmatic equivalent if they produce outputs that convey the same practical or meaningful information for the intended user, even if the outputs differ in format, structure, or additional details.
Semantic equivalence is a well understood concept in database community. In contrast to semantic equivalence, pragmatic equivalenceBaker (2018) is originated in natural language translation communities, focuses on maintaining the intended meaning or effect of an expression rather than its literal structure. For example, in natural language translation, a direct translation of the English idiom ’It rains cats and dogs’ into another language would fail to convey the intended meaning, as the phrase is nonsensical. Instead, the pragmatically equivalent translation would be ’rain heavily,’ which conveys the same idea of intense rainfall in a way that is meaningful to a non-English-speaking audience, without referencing cats or dogs. Similarly, in the context of SQL, pragmatic equivalence considers whether two queries achieve the same user intent or provide comparable practical information, even if their outputs or structures differ. This concept is particularly relevant in text-to-SQL systems, where multiple distinct yet valid queries can effectively fulfill the same user intent.
Based on the semantic definition of SQL equivalence (Definition 2.1), the Count bug is indeed a flaw for query engines and should be remedied immediately. On the other hand, we argue that based on Definition 2.2, which reflects the logical perspective and aligned with human perception, the two SQL queries in the Count Bug example are pragmatic equivalent. The fact that even experienced experts failed to identify this subtle difference indicates humans often recognize equivalence based on the more relaxed pragmatic notion. This observation motivates us to explore whether LLMs can serve as a more effective metric for evaluating text-to-SQL tasks, aligning better with human judgment.
2.2 Large Language Models Basics
Large language models (LLMs) are trained on extensive, web-scale datasets and leverage attention mechanisms Vaswani et al. (2017) combined with Reinforcement Learning with Human Feedback (RLHF) to generate high-quality, contextually relevant responses to input prompts. These models have demonstrated exceptional capabilities in solving complex tasks across diverse domains, such as natural language-to-SQL translation Pourreza and Rafiei (2023); Gao et al. (2023), data system optimizations Zhou et al. (2023), and data discovery Trummer (2023); Fernandez et al. (2023).
A defining characteristic of LLMs is their ability to function as few-shot learners Brown et al. (2020). One widely used approach to enhance their accuracy is in-context learning (ICL). Through ICL, LLMs utilize example demonstrations provided within the input prompt to improve task performance dynamically. Morevoer, when prompted with specific instructions and requirements, such as generating output in JSON format, LLMs adapt their responses to align with these constraints, ensuring accurate and tailored results.
3 Proposed Method
Motivated by recent methods Deng et al. (2023); Yao et al. (2023); Wei et al. (2022), we present our prompting techniques to obtain high quality judgments from Open-AI GPT models. Our prompting technique adheres to key aspects widely adopted in text-to-SQL works Li et al. (2023); Khatry et al. (2023); Liu et al. (2023); Taori et al. (2023): Clear Layout: We provide the two SQL queries that need to be compared at the beginning, then followed with the database schema, and end with the task (e.g., semantic equivalence or pragmatic equivalence) and instructions. Clear Context: We provide the database schema and explicitly included the primary and foreign key constraints, since they are crucial context information for deciding SQL equivalence. Prompt style: We adopted the code-representative approach, which has proven highly effective for database-related tasks Gao et al. (2023). An example prompt is illustrated below:
For evaluating semantic equivalence, we propose two distinct prompts. One approach utilizes a chain-of-thought (CoT) prompt, where the task and instructions explicitly guide the model to determine semantic equivalence by reasoning step-by-step. While the chain-of-thought prompt is generally effective, as shown in Figure 1, it often yields suboptimal results. This is particularly evident when language models merely scratch the surface, focusing on obvious similarities in the structure and logic of SQL queries rather than nuanced differences. This observation aligns with prior studies Wang et al. (2023); Zheng et al. (2023), which highlight that language models tend to overlook subtle distinctions when assessing equivalence and demonstrate limited reasoning capabilities in such contexts.
To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask LLMs to execute both queries on a simple database. If the outputs of the two SQL queries are identical, we then instruct LLMs to adjust the simple database (e.g., updating rows) and observe how the adjustments affect the outputs. If the modified database causes the two queries to produce different results, it indicates semantic inequivalence (i.e., a counterexample exists).
As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
In addition, we have added a specific instruction in both CoT and Miniature & Mull prompt: "The SQL predicate is case-sensitive when comparing string values". We have observed that LLMs often assume databases are case-insensitive. While many databases, such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013), are indeed case-insensitive by default, the Spider benchmark employs SQLite, which is case-sensitive by default. To address this, both the CoT and Miniature & Mull prompts explicitly specify that the database is case-sensitive when comparing string values.
For assessing pragmatic equivalence, we observed that LLMs often exhibit a tendency to label two SQL queries as equivalent based primarily on their structural and semantic similarity. This bias can lead to erroneous conclusions, as structural resemblance does not necessarily imply logical equivalence. To address this challenge, it is essential for LLMs to first identify and articulate the differences between the queries and then leverage this understanding to evaluate whether the queries are pragmatically equivalent. As a result, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent. This prompt follows the main ideas from the chain-of-thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries to understand their high-level objectives and then ask LLMs to determine if these two queries contain aligned logic or significant differences, as shown in Figure 1. This structured approach mitigates the LLMs’ inclination to focus solely on surface-level similarities. By requiring them to articulate reasoning steps, the prompt ensures a deeper evaluation of equivalence.
Moreover, we added specific instruction in Explain & Compare prompt regarding NULL values and case-sensitivity: "Database contains no NULL values". We have observed that LLMs often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that LLMs consider cases when NULL values can be inserted into a column with primary key constraint. In fact, the SQLite databases, by default, allow null values to be inserted to primary key column. However, in many database primary key constraints explicitly states that null values are not allowed Oracle . Although LLMs correctly consider that having NULL values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), these are not significant issue and does not affect the meaning of the query.
4 Evaluations
We evaluate the capabilities of popular large language models in identifying semantic and pragmatic equivalence between a pair of SQL queries. In particular, we aim to answer the following questions:
Is the Miniature & Mull prompt effective in accurately determining semantic equivalence?
Does LLMs using the Explain & Compare prompt achieve better alignment with human preferences than execution match?
We showcase the experimental results using the popular GPT-3.5-Turbo (gpt-3.5-turbo) and GPT-4o (gpt-4o) models. The models are accessed via OpenAI API.
Hyper-parameter: In the API call for all experiments, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 4096.
Few shot demonstrations. In addition to the zero-shot setting, we also investigate the in-context learning capabilities of LLMs in determining SQL equivalence. This is achieved by providing five static few-shot examples as part of the input context to guide the model’s reasoning process. For the semantic equivalence, the five static examples include three semantically equivalent SQL queries and two semantically inequivalent SQL queries. Similarly, for pragmatic equivalence, the five static examples consist of three equivalent and two inequivalent query pairs. The equivalent examples feature queries that share the same intent despite differences in their structure. For instance, queries like "SELECT birth_country FROM player" and "SELECT birth_city FROM player" are considered equivalent as both aim to determine the birth location, albeit at different levels of granularity.
4.1 Data Source
First, we compare our proposed Miniature & Mull prompt with the CoT prompt on i) Calcite test suites Begoli et al. (2018); Ding et al. (2023) and ii) Inequivalent Queries Gao et al. (2023). These two data sources include pairs of SQL queries that have been expertly labeled as semantically equivalent or non-equivalent.
Next, we present the results evaluating the alignment between our proposed LLM-SQL-Solver’s pragmatic equivalence judgments and human preferences. To conduct this assessment, we selected 70 pairs of SQL queries, comprising gold-standard SQL queries from the Spider benchmark Yu et al. (2018) and predicted SQL queries generated by LLMs Gao et al. (2023); Pourreza and Rafiei (2023); Liu et al. (2023). Each SQL pair was independently reviewed and labeled by at least three human experts proficient in SQL. The final label for each pair was determined through a majority voting process, ensuring consensus among the reviewers. This approach provides a robust benchmark for comparing the LLM’s pragmatic equivalence evaluations with human judgment. In addition to the human preference, we also provide the execution accuracy (EX) by executing these 70 queries over Spider databases. In all experiments, we use the open sourced EX evaluation code from GitHub ***https://github.com/THU-BPM/ChatGPT-sql.
Calcite test suites: It contains 232 pairs of equivalent queries (verified by expert) in the Calcite test suites.
Semantically Inequivalent Queries: It contains 180 pairs of semantic inequivalent queries. One of the query is the gold SQL in spider benchmark Yu et al. (2018) and the other is a predicted SQL query from DAIL-SQL Gao et al. (2023). Because they have different execution outputs, the databases in spider benchmark serve as counter examples.
Pragmatic Equivalence Queries: It contains 70 pairs of SQL queries with human preference label. Each pair has been independently reviewed and labeled by at least three human experts to capture human preference judgments.
4.2 Judging Semantic Equivalence
| Judge | Prompt | Semantic Eq. | Semantic Ineq. | ||
| 0-Shot | 5-Shot | 0-Shot | 5-Shot | ||
| ChatGPT | CoT | 53% | 52% | 63% | 46% |
| M&M | 85% (+32%) | 87% (+35%) | 67% (+4%) | 47% (+1%) | |
| GPT-4o | CoT | 81% | 83% | 85% | 83% |
| M&M | 84% (+3%) | 89% (+6%) | 86% (+1%) | 85% (+2%) | |
Recall that two queries cannot be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs. We used two different prompting methods CoT and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the CoT prompt. For determining the semantic equivalent queries (Calcite test suite), Miniature & Mull prompt using GPT-4o achieved around 90% accuracy. For determining the semantic inequivalent queries (DAIL-SQL results over Spider Benchmark), Miniature & Mull prompt using GPT-4o achieved around 86% accuracy. We can also observe that when using the weaker GPT-3.5-Turbo in judging semantically equivalent queries, Miniature & Mull outperform CoT prompt (asking LLMs to determine SQL equivalence and think step by step) by more than 30%. However, when judging the inequivalent SQL queries, we find that the GPT-3.5-Turbo significantly under-perform than the GPT-4o Turbo model.
Analyzing responses for identifying SQL-inequivalent queries, we observed a few reasons for GPT-3.5-Turbo’s low accuracy:
1.
Distinct entries vs duplicated entries. GPT-3.5-Turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated.
2.
Case-Sensitivity Ignored: GPT-3.5-Turbo often disregards the fact that the database is case-sensitive. Althought we have explicitly noted the database is string case-sensitive in the prompt, GPT-3.5-Turbo often fails to strictly adhere to the case-sensitivity requirement, resulting in inaccuracies when comparing queries. This observation aligns with findings in prior work Wen et al. (2024), which highlight GPT-3.5-Turbo’s challenges in reliably following instructions compared to stronger models such as GPT-4o.
3.
Overlooked Subtle Differences: GPT-3.5-Turbo may also fail to detect critical differences in column and table names. For instance, it might incorrectly treat distinct column names as equivalent or overlook the inclusion of additional columns or elements in a query. A specific example includes comparing AVG(capacity) with avg_capacity, where the former performs an average operation on the capacity column, while the latter refers to an entirely different column. Similarly, in cases like GROUP BY id versus GROUP BY student_id, GPT-3.5-Turbo may fail to recognize these distinctions, potentially leading to incorrect judgments.
4.
Hallucination. In the Miniature & Mull prompt, LLMs are instructed to execute both SQL queries on a simple database they create and then repeat the process on a modified version of the database to compare outputs. GPT-3.5-Turbo often returns incorrect results when executing SQL queries on its self-generated database instances. Moreover, even when the outputs are identical, it may mistakenly assert that the results differ. This hallucination and the resulting snowballing effect Zhang et al. (2023a) significantly contribute to GPT-3.5-Turbo’s poorer performance compared to GPT-4o.
When GPT-4o is used to determine SQL semantic equivalence, it consistently produces more accurate and reliable decisions for both equivalent and inequivalent queries. A notable observation is that GPT-4o reliably identifies the distinction between ORDER BY ASC/DESC LIMIT 1 and WHERE col = (SELECT MIN/MAX(col)). The former always returns a single row due to the LIMIT 1 clause, whereas the latter may return multiple rows if several share the same minimum or maximum value. This critical difference is often overlooked by GPT-3.5-Turbo. When using the CoT (Chain-of-Thought) prompt, GPT-4o achieves an accuracy of 82% – 85%, and with the Miniature & Mull prompt, it improves further to 85% – 90%.
Furthermore, we observe that, in general, adding few-shot demonstrations improves accuracy. However, for semantically inequivalent queries, incorporating the five static examples results in a significant drop in accuracy for both the CoT and Miniature & Mull prompts when using GPT-3.5-Turbo. The same few shot approach leads to only a slight accuracy decline when using GPT-4o. This discrepancy underscores the superior in-context learning capabilities of GPT-4o compared to GPT-3.5-Turbo. GPT-4o demonstrates a greater ability to interpret and generalize from the provided examples, enabling it to more effectively account for nuanced differences in semantic meaning. Conversely, GPT-3.5-Turbo struggles with leveraging the few-shot examples for semantically inequivalent queries, likely due to limitations in its contextual reasoning and pattern recognition abilities. This results in a diminished capacity to differentiate between queries with subtle semantic differences. Moreover, we believe that determining SQL semantic equivalence is an inherently challenging task for LLMs. It requires an understanding of both the syntax and the underlying logic of the queries. While few-shot learning provides valuable context, it may not always suffice to overcome the complexities involved, particularly for cases where the differences between queries are subtle or deeply tied to specific database schema (e.g., Primary Key and Foreign Key relationships).
Despite the aforementioned challenges in determining SQL semantic equivalence, our proposed Miniature & Mull prompt always achieve higher accuracy than using Cot prompt. This demonstrates the effectiveness of our approach in addressing some of the inherent difficulties of the task. As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.
To conclude this subsection, we also note that there are a few very rare cases that we observed in which the SQL queries are semantically equivalent but execution match determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by execution match but very rare.
| Judge | Data | Easy | Medium | Hard | Extra | Overall Acc | Rank |
| - | Spider Benchmark | 248 | 446 | 174 | 166 | - | - |
| EX | Chat-SQL | 221/248 | 331/446 | 101/174 | 72/166 | 70.1% | 3 |
| DIN-SQL | 228/248 | 382/446 | 132/174 | 101/166 | 81.5% | 2 | |
| DAIL-SQL | 226/248 | 392/446 | 132/174 | 104/166 | 82.6% | 1 | |
| ChatGPT | Chat-SQL | 223/248 | 397/446 | 149/174 | 128/166 | 86.8% | 3 |
| DIN-SQL | 228/248 | 391/446 | 153/174 | 133/166 | 87.5% | 2 | |
| DAIL-SQL | 235/248 | 407/446 | 158/174 | 142/166 | 91.1% | 1 | |
| GPT-4o | Chat-SQL | 226/248 | 328/446 | 129/174 | 104/166 | 76.1% | 3 |
| DIN-SQL | 231/248 | 376/446 | 150/174 | 126/166 | 85.4% | 2 | |
| DAIL-SQL | 228/248 | 397/446 | 153/174 | 129/166 | 87.7% | 1 |
4.3 Judging Text-to-SQL Generation
In this subsection, we investigate whether large language models (LLMs) can effectively assess the quality of text-to-SQL generations. Specifically, we compare our proposed Explain & Compare method with the widely used execution match approach. Additionally, we examine which evaluation metric (i.e., LLM judgments or execution match) aligns more closely with user opinions, providing deeper insights into the effectiveness of these approaches.
We first introduce three state-of-the art text-to-SQL algorithms for the Spider benchmark.
DIN-SQL Pourreza and Rafiei (2023) ranks the fourth in the leader board and it decompose the text-to-SQL translation problem into sub tasks such as schema linking and classifying the level of difficulties. DIN-SQL uses the GPT-4 model with few-shot prompting.
Chat-SQL Huang et al. (2023) directly generate queries using GPT-3.5-Turbo with the zero-shot basic prompt.
In Table 2, we present the breakdown of the alignment between human preference and the execution match and the LLM judgment using Explain & Compare prompt on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL).
First, we observe that the relative ranking among Chat-SQL, DIN-SQL, and DAIL-SQL remains consistent when transitioning from execution match accuracy to LLM-based evaluations. Given that Chat-SQL relies on SQL queries generated by the weaker model, GPT-3.5-Turbo, using zero-shot prompting, it is expected that the SQL queries produced by DIN-SQL and DAIL-SQL would demonstrate higher accuracy. Furthermore, DAIL-SQL stands out by extensively exploring various prompting techniques for text-to-SQL translation, employing the most effective combination of strategies, and incorporating resource-intensive self-consistency methods. As a result, DAIL-SQL is more likely to generate the most accurate responses. As shown in Table 2, the ranking among Chat-SQL, DIN-SQL, and DAIL-SQL based on execution accuracy, GPT-3.5-Turbo judgments, and GPT-4o judgments with our proposed Explain & Compare, all aligned with our expectation.
Additionally, when we examine the accuracy across different difficulty levels (i.e., easy, medium, hard, and extra) in the Spider benchmark, it becomes evident that the primary discrepancies between the execution metric and the LLMs’ judgment metric arise from questions classified as hard and extra difficulty. LLMs’ judgments, based on the pragmatic equivalence, are less strict than the execution match metric. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). LLMs using Explain & Compare prompt are able to identify the pragmatic equivalence between these two queries (i.e., identifying the top/bottom results based on col) despite their syntactic differences. However, since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent depending on the underlying database (i.e., whether there are duplicate records sharing the same maximum or minimal value).
| Judge | Data | Agreement |
| GPT-3.5-Turbo | Chat-SQL | 75.8% |
| DIN-SQL | 83.4% | |
| DAIL-SQL | 85.5% | |
| GPT-4o | Chat-SQL | 82.6% |
| DIN-SQL | 87.2% | |
| DAIL-SQL | 89.5% |
4.3.1 Agreement between LLMs’ Judgments and Execution Accuracy
For the majority of questions, the judgments made by large language models (LLMs) align well with execution-based (EX) accuracy. Specifically, as shown in Table 3, the agreement rates between LLMs’ judgments (based on pragmatic equivalence) and execution accuracy range from 76% to 90%. This strong alignment is an encouraging signal, reflecting both the quality of the Spider benchmark and the reliability of LLMs’ judgments. The Spider benchmark is renowned for its high-quality databases, developed with significant human effort—approximately 1,000 hours invested in their creation Yu et al. (2018). This meticulous design ensures the databases are robust, enabling execution-based metrics to produce reliable results. The benchmark’s reputation has been widely celebrated within the community for setting a standard in text-to-SQL evaluation.
However, creating such high-quality databases is resource-intensive and does not easily scale, limiting the broader applicability of execution-based metrics. In this context, the high agreement between execution accuracy and LLM-based judgments is particularly desirable. It suggests that LLMs, when guided by pragmatic equivalence, can provide a scalable and reliable alternative for evaluating query accuracy, reducing the dependence on expensive database creation while maintaining strong evaluation standards.
4.3.2 Alignment with Human Preference
While LLM judgments and execution-based metrics agree on the majority of questions, there remain many instances where their evaluations diverge. In this subsection, we analyze which metric, LLM judgments or execution accuracy, better aligns with human preferences. By examining these discrepancies, we aim to uncover the strengths and weaknesses of each evaluation approach in accurately assessing the quality of text-to-SQL systems.
| Judge | Few Shot | Human Prefer |
| Execution Match | - | 40.0% |
| GPT-3.5-Turbo | 0 Shot | 72.9% |
| 5 Shot | 74.3% | |
| GPT-4o | 0 Shot | 68.6% |
| 5 Shot | 78.6% |
Human Preferences To evaluate the pragmatic equivalence between two SQL queries in text-to-SQL service, we use the human preference approach. Higher alignment to human preference indicates better judgment. At least three experts in SQL have manually checked the equivalence of the 70 selected queries from the state-of-the-art text-to-SQL algorithms Pourreza and Rafiei (2023); Gao et al. (2023); Huang et al. (2023). These 70 questions include 9 easy questions, 27 medium questions, 14 hard questions, and 20 extra-hard questions. They were randomly selected from cases where an LLM’s judgment disagreed with the execution match metric. To ensure the quality of human judgment, the annotations are conducted independently, and we use the majority vote to determine the final label for each pair of SQL queries.
In Table 4, we find that execution match only aligned 40% of the time with human preferences. When using the Explain & Compare prompt for determining the pragmatic equivalence, LLMs are consistently better aligned with human preference than execution match, as shown in Table 4. The pragmatic notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning and intent, mimicking the opinions of text-to-SQL users. By reasoning into the query logic and determine whether the two SQL queries share the same logic and have the same intent, LLMs’ judgments become better aligned with human preference than execution accuracy. Moreover, when five static examples are included in the prompt, GPT-3.5-Turbo aligned with human preference 74.3% and GPT-4o aligned with human preference 78.6%. The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
Diving deeper into the questions where LLMs failed to align with human preferences, we find that these questions often lack unanimous agreement among experts, as the definition of pragmatic equivalence can be highly subjective. For instance, consider the following two SQL queries:
The queries differ in their filtering criteria: the first uses the airport name to filter for ’Alton’, while the second uses the airport code to filter for ’Alton’. Among experts, two considered these queries pragmatically inequivalent, as the filtering criteria are based on different attributes, which could yield totally different output. However, one expert judged them to be pragmatically equivalent, as they have the same intent of finding location based on predicate ’Alton’. LLMs, failing to converge to the majority opinion, interpreted these queries as having the same meaning, possibly oversimplifying the nuanced distinctions in their semantic intent. An exciting avenue for future research is to develop customizable text-to-SQL evaluation metrics tailored to individual user preferences, leveraging in-context learning to adapt the system’s understanding and judgments based on specific user requirements or domain contexts.
5 Discussion and Future Works
In this section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver. Moreover, Concurrent with our research, the chain-of-table approach Wang et al. (2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt. However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
5.1 Applications of LLM-SQL-Solver
First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases. Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub-optimal performance. We believe that during the SQL rewrite phase of the query optimization engine, the LLM-SQL-Solver can be directly used as a tool to help checking semantic equivalence, ensuring bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur and also explore more SQL rewrite possibilities which may lead to better query performance.
For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from a user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents). Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al. (2018). Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database. This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information. As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases (save costs) and without accessing any stored data (ensure data privacy). Based on our extensive experiment evaluations, we believe that LLM-SQL-Solver (with pragmatic notion) can be used as a better metric for text-to-SQL benchmarks.
5.2 Future Plans
There are areas of improvement to this work. First, we find LLMs may hallucinate in executing the queries over simple databases. By equipping LLMs with tools such as Python pandas library, LLMs can reply on these tool for observing the execution results of SQL queries Zhuang et al. (2023). By reducing the hullicination, we believe that our proposed Miniature & Mull prompt will have an increased accuracy in determining SQL semantic equivalence.
In addition, our evaluations of pragmatic equivalence were conducted on a set of 70 questions, requiring experts to manually label the equivalence of each pair of SQL queries. This process is time-consuming, as even experienced SQL experts need approximately 10 minutes to thoroughly analyze and determine the equivalence of a single query pair. Despite the relatively small dataset size, we observed consistent trends in the alignment between LLM judgments and execution matches. Notably, GPT-4o’s responses consistently demonstrated better alignment with human preferences compared to other models, showcasing its ability to better capture nuanced distinctions in query intent. Expanding the dataset would allow for a more robust analysis and further validation of these trends, providing deeper insights into the strengths and weaknesses of current models.
Moreover, our analysis revealed that pragmatic equivalence is often highly subjective, depending on the context and the user’s interpretation of query intent. This subjectivity underscores the need for designing customizable evaluation metrics tailored to individual users. Such personalized metrics could adapt to specific user needs and preferences, ultimately improving the quality and relevance of text-to-SQL systems for diverse applications. Future research in this direction could significantly enhance the usability and practicality of these systems, enabling them to cater to a broader range of use cases and user requirements.
6 Conclusion
In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT-4o), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with pragmatic equivalence notion, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, using our proposed LLM-SQL-Solver as the evaluation metric for text-to-SQL translation ensures data privacy by eliminating the need to access stored user data, and it reduces costs by removing the requirement to construct sophisticated databases, making the evaluation process more efficient and scalable.
Acknowledgment
We thank Tom Shangguan for his assistance in organizing and managing the data files used in this research.